An organic compound shows two peaks at m/e 64 and 66 (3 : 1 ratio). Other peaks are obtained
at m/e 49, 51 (3 : 1) and 29. Identify the compound.
1
Expert's answer
2019-11-04T06:23:57-0500

Answer on Question #97911 – Chemistry – Other


Task:

An organic compound shows two peaks at m/e 64 and 66 (3:1 ratio). Other peaks are obtained at m/e 49, 51 (3:1) and 29. Identify the compound.


Solution:

Such look of mass spectrum eloquently indicates the presence of chlorine. In nature, there are two stable isotopes 35Cl and 37Cl with a percentage distribution of 75.8% and 24.2%. In this case, a compound containing chlorine would have two peaks separated by 2 units and a peak containing a chlorine isotope 35Cl would be about 3 times higher than the peak at M+2, which is observed in our case.

The presence of a peak at m/e = 29 indicates the ethyl group CH3-CH2-.

Peaks at m/e = 49 and 51 clearly occur from the separation of only the methyl group CH3- from the compound and may correspond to the formation of chloromethyl radical -CH2Cl.

Then the peak at m/e = 64 is Ethyl chloride - CH3-CH2-Cl. 


Answer: the compound = ethyl chloride = CH3-CH2-Cl.